OpenStudy (anonymous):

If sin A = 2/3 where 0° < A < 90°, what is the value of sin 2A

OpenStudy (anonymous):

1) (2radical5)/3 2)(2radical5)/9 3)(4radical5)/9 4)-(4radical5)/9

OpenStudy (danjs):

2*sin(A)*cos(A)

OpenStudy (danjs):

double angle, what is cos(A) if the sin(A)=2/3? It is the point (x,y) = (cos(A), Sin(A)) on the unit circle

OpenStudy (danjs):

Recall: \[\sin^2(A) + \cos^2(A) = 1\]

OpenStudy (danjs):

\[Sin(2A) = 2*Sin(A)*Cos(A)\]

OpenStudy (anonymous):

i know that but i am having trouble using this to help me find the correct answer

OpenStudy (danjs):

If you want to find the cosine first, use \[(2/3)^2 + \cos^2(A) = 1\] \[\cos(A) = \sqrt{1-4/9} = \sqrt{5/9} = \frac{ \sqrt{5} }{ 3 }\]

OpenStudy (danjs):

Then From double angle \[Sin(2A) = 2*Sin(A)*Cos(A) = 2*\frac{ 2 }{ 3 }*\frac{ \sqrt{5} }{ 3 }\]

OpenStudy (anonymous):

(4radical5)/9?